Batteries are one of the major imports of the US from China. Batteries cost 10 Yuan in China.

a) If the exchange rate is 7 Yuan/$, how much does the battery cost in the US?

b) If the Dollar appreciates to 8 Yuan/$, how much does the battery cost in the US now? Assume that the price of the battery does not change.

c) What happens to the import of batteries from China? What happens to the NX of the US?

d) If the Yuan appreciates to 6 Yuan/$, how much does the battery cost in the US now? Assume that the price of the battery does not change.

e) What happens to the import of batteries from China? What happens to the NX of the US?
Answer #1

(a)

Cost of battery in US = 10/7 = $1.43

(b)

New Cost of battery in US = 10/8 = $1.25

(c)

Since cost of battery in US in terms of dollar has decreased, US will increase its impot demand from China. So, import will increase from China, and net exports will decrease in US.

(d)

New Cost of battery in US = 10/6 = $1.67

(e)

Since cost of battery in US in terms of dollar has increased, US will decrease its impot demand from China. So, import will decrease from China, and net exports will increase in US.